I need help,I am trying to help my daughter do her algebra homework and do not remember the formula for these word problems and she is in need of help.
Larry is 15 Years oldr than Beth, In 4 Years larry will be one year older than Twice beths age. Find their ages today.
Thank you for your help.

Answer by ankor@dixie-net.com(12695)

let L = Larry's present age
Let B = Beth's present age
;
Write an equation for each statement:
:
"Larry is 15 Years older than Beth,"
L = B + 15
:
"In 4 Years Larry will be one year older than Twice Beth's age."
L + 4 = 2(B+4) + 1
Multiply what's inside the brackets, Simplify
L + 4 = 2B + 8 + 1
L + 4 = 2B + 9
Subtract 4 from both sides
L = 2B + 9 - 4
L = 2B + 5
:
Find their ages today.
Substitute (B+15) for L in the above equation
B + 15 = 2B + 5
Combine like terms
15 - 5 = 2B - B
10 = B
Beth is 10 yrs old
then
15 + 10 = 25 yrs is Larry's present age
:
:
Check these solution in the statement:
""In 4 Years Larry will be one year older than Twice Beth's age."
25 + 4 = 2(10+4) + 1
29 = 2(14) + 1; confirms our solution